### v1

@@ Abstract
Context. TOI-396 is an F6 V, bright, naked-eye star ( 6.4) orbited by three transiting small ( 2 ) planets discovered thanks to space-based photometry from 2 TESS sectors. The orbital periods of the two innermost planets, namely TOI-396 b and c, are close to the 5:3 commensurability ( 3.6 d and 6.0 d), suggesting that the planets might be trapped in a mean motion resonance (MMR).
Aims. To measure the masses of the three planets, refine their radii, and investigate whether planets b and c are in MMR, we carried out HARPS radial velocity (RV) observations of TOI-396 and retrieved archival high-precision transit photometry from 4 TESS sectors.
Methods. We extracted the RVs via a skew-normal fit onto the HARPS cross correlation functions and performed a Markov chain Monte Carlo joint analysis of the Doppler measurements and transit photometry, while employing the breakpoint method to remove stellar activity from the RV time series. We also performed a transit timing variation (TTV) dynamical analysis of the system and simulated the temporal evolution of the TTV amplitudes of the three planets following an N-body numerical integration.
Results. Our analysis confirms that the three planets have similar sizes (; ; ), in agreement with previous findings. Notably, our measurements are 1.4 times more precise, thanks to the use of two additional TESS sectors. For the first time, we determine RV masses for TOI-396 b and d, and find them to be and , implying bulk densities of g cm-3 and g cm-3, respectively. Our results suggest a quite unusual system architecture, with the outermost planet being the densest. Based on a frequency analysis of the HARPS activity indicators and TESS light curves, we find the rotation period of the star to be d, in agreement with the value predicted from -based empirical relations. The Doppler reflex motion induced by TOI-396 c remains undetected in our RV time series, likely due to the proximity of the planet’s orbital period to the star’s rotation period. We also discovered that TOI-396 b and c display significant TTVs. While the TTV dynamical analysis returns a formally precise mass for TOI-396 c of , the result might not be accurate, owing to the poor sampling of the TTV phase. We also conclude that TOI-396 b and c are close to-, but out of- the 5:3 MMR.
@@
If planets are found in mean motion resonance (MMR; e.g. Lee & Peale, 2002; Correia et al., 2018), systems can also shed light on the migration mechanisms during formation, as well as on the impact of tidal effects occurring later on (e.g. Delisle et al., 2012; Izidoro et al., 2017). In addition, planets in, or close to, MMR likely exhibit transit timing variations (TTVs; see e.g. Agol et al., 2005; Agol & Fabrycky, 2018; Leleu et al., 2021) that enable one to infer the planetary masses without necessarily relying on radial velocity (RV) measurements, which are not always possible (e.g. Hatzes, 2016).
@@
TOI-396 represents an interesting laboratory to test these theories as it is the brightest star known so far to host three transiting planets (Vanderburg et al., 2019) after Lupi (Delrez et al., 2021). After analysing two sectors from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015), Vanderburg et al. (2019) found that the three planets have radii of 2 and orbital periods of 3.6, 6.0, and 11.2 d, with TOI-396 c and b showing a period commensurability close to the 5:3 ratio. Following the notation introduced in Mishra et al. (2023), the three planets are similar in terms of radii, and one may wonder whether this architecture class is kept also in the mass-period diagram. Mishra et al. (2023) found a positive and strong correlation of the coefficients of similarity between radii and masses, though exceptions are possible (e.g. Weiss & Marcy, 2014; Otegi et al., 2020a, 2022).
In this work we complement the photometric analysis of new TESS light curves (LCs) with RV observations taken with the High Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al., 2003) spectrograph to refine the planet radii and constrain for the first time the planetary masses. Considering the possible 5:3 MMR between TOI-396 c and b, we also simulate the evolution in time of the TTV amplitudes. This paper is organised as follows. Section 2 presents the stellar properties and Section 3 describes the photometric and RV data. After outlining the method to jointly analyse the TESS LCs and the HARPS RV time series in Section 4, we present the corresponding results in Section 5. We attempt to dynamically model TTV and RV data simultaneously and we track the temporal evolution of the TTV signals in Section 6, study the planets’ internal structure in Section 7, and explore the prospects for characterising the system with the James Webb Space Telescope (JWST; Gardner et al., 2006) in Section 8. Finally Section 9 gathers the conclusions.
@@ 2 Host star characterisation
To double-check the derived spectroscopy parameters we performed an additional analysis employing ARES+MOOG (Sousa et al., 2021; Sousa, 2014; Santos et al., 2013). In detail, we used the latest version of ARES 222The last version, ARES v2, can be downloaded at https://github.com/sousasag/ARES. (Sousa et al., 2007, 2015) to consistently measure the equivalent widths (EW) for the list of iron lines presented in Sousa et al. (2008). Following a minimisation process, we then find the ionization and excitation equilibria to converge for the best set of spectroscopic parameters. This process uses a grid of Kurucz model atmospheres (Kurucz, 1993) and the radiative transfer code MOOG (Sneden, 1973). We obtained K, dex, , and km s-1. In this process we also derived a more accurate trigonometric surface gravity () using recent GAIA data following the same procedure as described in Sousa et al. (2021). In the end ARES+MOOG provides consistent values when compared with the ones derived with SME.
@@
The stellar radius , mass , and age were derived homogeneously using the isochrone placement algorithm (Bonfanti et al., 2015, 2016) and its capability of interpolating a flexible set of input parameters within pre-computed grids of PARSEC333http://stev.oapd.inaf.it/cgi-bin/cmd. v1.2S (Marigo et al., 2017) isochrones and tracks. For each magnitude listed in Table 1, we performed an isochrone placement run by inserting the spectroscopic parameters derived above, the Gaia parallax (Gaia Collaboration et al., 2023, offset-corrected following Lindegren et al. 2021), and the magnitude value to obtain an estimate for the stellar radius, mass, and age along with their uncertainties. From these results, we built the corresponding Gaussian probability density functions (PDFs) and then we merged (i.e. we summed) the PDFs derived from the different runs to obtain robust estimates for , , and . The radius derives essentially from , , and the stellar magnitude, while and are much more model-dependent, therefore we conservatively inflate their internal uncertainties by and , respectively, following Bonfanti et al. (2021). Our adopted stellar parameters are listed in Table 1.
@@ 3.1 TESS photometry
We analyse all available TESS time series, including the Sector 3 and 4 data already presented in (Vanderburg et al., 2019). For Sector 3 we used the TESS Asteroseismic Science Operation Center (TASOC) photometry (Handberg et al., 2021; Lund et al., 2021), while for the other sectors we analysed the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) light curves generated by the TESS Science Processing Operation Center (SPOC) pipeline Jenkins et al. (2016), which removes the majority of instrumental artefacts and systematic trends (Smith et al., 2012; Stumpe et al., 2012, 2014). We rejected those data marked with a bad quality flag and we performed a 5 median-absolute-deviation (MAD) clipping of flux data points. After that, we built our custom light curves (LCs) by splitting the TESS sectors in temporal windows centred around the transit events keeping 4 h of out-of-transit data points both before and after the transit for detrending purposes. We ended up with 41 TESS LCs reporting the epoch of observation (t), the flux and its error, and further ancillary vectors available from the TESS science data product, that is mom_centr1, mom_centr2 (hereafter denoted with x and y, respectively), pos_corr1, and pos_corr2 (hereafter denoted with pc1 and pc2, respectively)555Please refer to https://tasoc.dk/docs/EXP-TESS-ARC-ICD-TM-0014-Rev-F.pdf for further details about the output of the TESS science data products..
@@ 3.2 HARPS high-resolution spectroscopy
We performed the radial velocity (RV) follow-up of TOI-396 with the HARPS spectrograph mounted at the ESO-3.6 m telescope at La Silla Observatory (Chile). We acquired 77 high resolution ( 115 000) spectra between 31 January and 27 July 2019 (UT), covering a baseline of 177 days, as part of the follow-up programs of TESS systems carried out with the HARPS spectrograph (IDs: 1102.C-0923, 1102.C-0249, 0102.C-0584; PIs: Gandolfi, Armstrong, De Medeiros). One additional spectrum was acquired during a technical night (ID: 60.A-9700) in February 2019. Following Dumusque et al. (2011b), we averaged out p-mode stellar pulsations by setting the exposure time to 900 s, which led to a median signal-to-noise (S/N) ratio of 320 per pixel at 550 nm. We used the second fibre of the HARPS spectrograph to simultaneously observe a Fabry-Perot lamp and trace possible instrumental drift down to the sub-m s-1 level (Wildi et al., 2010). We reduced the data using the dedicated HARPS data reduction software (DRS; Pepe et al., 2002; Lovis, C. & Pepe, F., 2007) and computed the cross-correlation function (CCF) for each spectrum using a G2 numerical mask (Baranne et al., 1996).
We then performed a Skew-Normal (SN) fit onto each CCF (Simola et al., 2019) to extract the stellar radial velocity along with its error, the full-width-at-half-maximum (FWHMSN), the contrast (), and the skewness parameter () of the CCF. The advantages of using an SN-fit rather than a Normal fit are thoroughly discussed in Simola et al. (2019), while the SN-fitting details are outlined in, e.g., Bonfanti et al. (2023); Luque et al. (2023); Fridlund et al. (2024). After the SN-based extraction, we ended up with an RV time series, whose ancillary vectors (FWHMSN, , ) are the activity indicators used to constrain the polynomial basis to model and detrend the RV component of the stellar activity (see Table LABEL:tab:RVdata).
As stellar activity is not stationary, the correlations between the RV observations and the activity indicators are expected to change over time as discussed in Simola et al. (2022), who proposed to apply the breakpoint (bp) technique (Bai & Perron, 2003) to check whether these correlation changes are statistically significant. If so, the bp algorithm finds the optimal locations of correlation changes, which defines the segmentation characterised by the lowest Bayesian Information Criterion (BIC; Schwarz, 1978). Indeed, we found one breakpoint at observation 48 (BJD = 2458667.940719; with respect to the zero-breakpoint solution). Thus, we split our RV time series in two piecewise stationary segments and detrended it on a chunk-wise base, rather than performing a global detrending over the whole time series, similarly to what has already been done in Bonfanti et al. (2023); Luque et al. (2023).
@@ 4 Methods
We jointly analysed the TESS LCs and the RV time series within a Markov chain Monte Carlo (MCMC) framework using the MCMCI code (Bonfanti & Gillon, 2020). When fitting the TESS LCs extracted from Sector 3 that has a cadence of 30 min, we generated the transit model using a cadence of 2 min (the same as the other TESS sectors) and then rebinned it to 30 min following Kipping (2010), who warns that long-cadence photometry may lead to retrieve erroneous system parameters. We imposed Normal priors on the input stellar parameters (i.e. , [Fe/H], , and ), as derived in Section 2 with a twofold aim: (i) the induced prior on the mean stellar density (via and ) helps the convergence of the transit model; (ii) limb darkening (LD) coefficients for the TESS filter may be retrieved following interpolation within Atlas9-based666http://kurucz.harvard.edu/grids.html. grids that were pre-computed using the get_lds.py code777https://github.com/nespinoza/limb-darkening by Espinoza & Jordán (2015). We then set Normal priors on the quadratic LD coefficients using the values coming from the grid interpolation (i.e. and ) after re-parameterising them following Holman et al. (2006). On the planetary side, we adopted unbounded (except for the physical limits) uniform priors on the transit depth d, the impact parameter , the orbital period , the transit timing , and the RV semi-amplitude , while we set the eccentricity and the argument of periastron for all planets. We come back to the assumption on the eccentricity below. The specific parameterisations of the jump parameters (aka step parameters) are outlined in Bonfanti & Gillon (2020).
@@
The LCs and the RV time series were detrended simultaneously during the MCMC analysis using polynomials. To assess the polynomial orders to be associated to the different detrending parameters for each time series, we first launch several MCMC preliminary runs made of 10 000 steps where we varied only one polynomial order at a time. We then selected the best detrending baseline (see Table 7) as the one having the lowest BIC. After that, we performed a preliminary MCMCI run comprising 200 000 steps to evaluate the contribution of both the white and red noise in the LCs following Pont et al. (2006); Bonfanti & Gillon (2020), so to properly rescale the photometric errors and get reliable uncertainties on the output parameters. Finally, three independent MCMCI runs comprising 200 000 steps each (burn-in length equal to 20%) were performed to build the posterior distributions of the output parameters after checking their convergence via the Gelman-Rubin statistic (; Gelman & Rubin, 1992).
@@
We also tested the possibility of eccentric orbits by imposing uniform priors on (, ) either bounded to imply or completely unbounded (except for the physical limits). The wider the eccentricity range to be explored by the MCMC scheme, the poorer the parameter convergence, which suggests that the available data are not enough to well constrain the planetary eccentricities. Moreover, the MCMCI runs with are disfavoured by the BIC criterion (e.g. Kass & Raftery, 1995; Trotta, 2007) as well, in fact we obtained . This is also in agreement with the simulations performed by Vanderburg et al. (2019), who suggested that the periods’ commensurability state of planets b and c is more likely maintained if the system is characterised by low eccentricities. Therefore, we adopted the circular solution as the reference one.
5 LC & RV data analysis results
5.1 Joint LC & RV analysis with linear ephemerides

### v2

@@ Abstract
Context. TOI-396 is an F6 V bright naked-eye star ( 6.4) orbited by three small ( 2 ) transiting planets discovered thanks to space-based photometry from two TESS sectors. The orbital periods of the two innermost planets, namely TOI-396 b and c, are close to the 5:3 commensurability ( 3.6 d and 6.0 d), suggesting that the planets might be trapped in a mean motion resonance (MMR).
Aims. To measure the masses of the three planets, refine their radii, and investigate whether planets b and c are in MMR, we carried out HARPS radial velocity (RV) observations of TOI-396 and retrieved archival high-precision transit photometry from four TESS sectors.
Methods. We extracted the RVs via a skew-normal fit onto the HARPS cross-correlation functions and performed a Markov chain Monte Carlo joint analysis of the Doppler measurements and transit photometry while employing the breakpoint method to remove stellar activity from the RV time series. We also performed a transit timing variation (TTV) dynamical analysis of the system and simulated the temporal evolution of the TTV amplitudes of the three planets following an N-body numerical integration.
Results. Our analysis confirms that the three planets have similar sizes (; ; ) and is thus in agreement with previous findings. However, our measurements are 1.4 times more precise thanks to the use of two additional TESS sectors. For the first time, we have determined the RV masses for TOI-396 b and d, finding them to be and , which implies bulk densities of g cm-3 and g cm-3, respectively. Our results suggest a quite unusual system architecture, with the outermost planet being the densest. Based on a frequency analysis of the HARPS activity indicators and TESS light curves, we find the rotation period of the star to be d, in agreement with the value predicted from -based empirical relations. The Doppler reflex motion induced by TOI-396 c remains undetected in our RV time series, likely due to the proximity of the planet’s orbital period to the star’s rotation period. We also discovered that TOI-396 b and c display significant TTVs. While the TTV dynamical analysis returns a formally precise mass for TOI-396 c of , the result might not be accurate, owing to the poor sampling of the TTV phase. We also conclude that TOI-396 b and c are close to but out of the 5:3 MMR.
@@
If planets are found in mean motion resonance (MMR; e.g. Lee & Peale, 2002; Correia et al., 2018), systems can also shed light on the migration mechanisms during formation, as well as on the impact of tidal effects occurring later on (e.g. Delisle et al., 2012; Izidoro et al., 2017). In addition, planets in, or close to, MMR likely exhibit transit timing variations (TTVs; see e.g. Agol et al., 2005; Agol & Fabrycky, 2018; Leleu et al., 2021) that enable one to infer planetary masses without necessarily relying on radial velocity (RV) measurements, which are not always possible (e.g. Hatzes, 2016).
@@
The object TOI-396 represents an interesting laboratory to test these theories as it is the brightest star known so far to host three transiting planets (Vanderburg et al., 2019) after Lupi (Delrez et al., 2021). After analysing two sectors from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015), Vanderburg et al. (2019) found that the three planets have radii of 2 and orbital periods of 3.6, 6.0, and 11.2 d, with TOI-396 c and b showing a period commensurability close to the 5:3 ratio. Following the notation introduced in Mishra et al. (2023), the three planets are ‘similar’ in terms of radii, and one may wonder whether this architecture class is kept also in the mass-period diagram. Mishra et al. (2023) found a positive and strong correlation of the coefficients of similarity between radii and masses, though exceptions are possible (e.g. Weiss & Marcy, 2014; Otegi et al., 2020a, 2022).
In this work, we complement the photometric analysis of new TESS light curves (LCs) with RV observations taken with the High Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al., 2003) spectrograph to refine the planet radii and constrain for the first time the planetary masses. Considering the possible 5:3 MMR between TOI-396 c and b, we also simulate the evolution in time of the TTV amplitudes. This paper is organised as follows: Section 2 presents the stellar properties, and Sect. 3 describes the photometric and RV data. After outlining the method to jointly analyse the TESS LCs and the HARPS RV time series in Sect. 4, we present the corresponding results in Sect. 5. We attempt to dynamically model TTV and RV data simultaneously and track the temporal evolution of the TTV signals in Sect. 6, and we study the planets’ internal structure in Sect. 7 and explore the prospects for characterising the system with the James Webb Space Telescope (JWST; Gardner et al., 2006) in Sect. 8. Finally Sect. 9 gathers the conclusions.
@@ 2 Host star characterisation
To double-check the derived spectroscopy parameters we performed an additional analysis employing ARES+MOOG (Sousa et al., 2021; Sousa, 2014; Santos et al., 2013). In detail, we used the latest version of ARES 222The last version, ARES v2, can be downloaded at https://github.com/sousasag/ARES. (Sousa et al., 2007, 2015) to consistently measure the equivalent widths (EW) for the list of iron lines presented in Sousa et al. (2008). Following a minimisation process, we then find the ionisation and excitation equilibria to converge for the best set of spectroscopic parameters. This process uses a grid of Kurucz model atmospheres (Kurucz, 1993) and the radiative transfer code MOOG (Sneden, 1973). We obtained K, dex, , and km s-1. In this process we also derived a more accurate trigonometric surface gravity () using recent GAIA data following the same procedure as described in Sousa et al. (2021). In the end ARES+MOOG provides consistent values when compared with the ones derived with SME.
@@
The stellar radius , mass , and age were derived homogeneously using the isochrone placement algorithm (Bonfanti et al., 2015, 2016) and its capability of interpolating a flexible set of input parameters within pre-computed grids of PARSEC333http://stev.oapd.inaf.it/cgi-bin/cmd. v1.2S (Marigo et al., 2017) isochrones and tracks. For each magnitude listed in Table 1, we performed an isochrone placement run by inserting the spectroscopic parameters derived above, the Gaia parallax (Gaia Collaboration et al., 2023, offset-corrected following Lindegren et al. 2021), and the magnitude value to obtain an estimate for the stellar radius, mass, and age along with their uncertainties. From these results, we built the corresponding Gaussian probability density functions (PDFs) and then we merged (i.e. we summed) the PDFs derived from the different runs to obtain robust estimates for , , and . The radius derives essentially from , , and the stellar magnitude, while and are much more model-dependent; therefore we conservatively inflated their internal uncertainties by and , respectively, following Bonfanti et al. (2021). Our adopted stellar parameters are listed in Table 1.
@@ 3.1 TESS photometry
We analyse all available TESS time series, including the Sector 3 and 4 data already presented in (Vanderburg et al., 2019). For Sector 3 we used the TESS Asteroseismic Science Operation Center (TASOC) photometry (Handberg et al., 2021; Lund et al., 2021), while for the other sectors we analysed the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) LCs generated by the TESS Science Processing Operation Center (SPOC) pipeline Jenkins et al. (2016), which removes the majority of instrumental artefacts and systematic trends (Smith et al., 2012; Stumpe et al., 2012, 2014). We rejected those data marked with a bad quality flag and we performed a 5 median-absolute-deviation (MAD) clipping of flux data points. After that, we built our custom LCs by splitting the TESS sectors in temporal windows centred around the transit events keeping 4 h of out-of-transit data points both before and after the transit for de-trending purposes. We ended up with 41 TESS LCs reporting the epoch of observation (t), the flux and its error, and further ancillary vectors available from the TESS science data product, that is mom_centr1, mom_centr2 (hereafter denoted with x and y, respectively), pos_corr1, and pos_corr2 (hereafter denoted with pc1 and pc2, respectively)555Please refer to https://tasoc.dk/docs/EXP-TESS-ARC-ICD-TM-0014-Rev-F.pdf for further details about the output of the TESS science data products..
@@ 3.2 HARPS high-resolution spectroscopy
We performed the radial velocity (RV) follow-up of TOI-396 with the HARPS spectrograph mounted at the ESO-3.6 m telescope at La Silla Observatory (Chile). We acquired 77 high resolution ( 115 000) spectra between 31 January and 27 July 2019 (UT), covering a baseline of 177 days, as part of the follow-up programs of TESS systems carried out with the HARPS spectrograph (IDs: 1102.C-0923, 1102.C-0249, 0102.C-0584; PIs: Gandolfi, Armstrong, De Medeiros). One additional spectrum was acquired during a technical night (ID: 60.A-9700) in February 2019. Following Dumusque et al. (2011b), we averaged out p-mode stellar pulsations by setting the exposure time to 900 s, which led to a median signal-to-noise (S/N) ratio of 320 per pixel at 550 nm. We used the second fibre of the HARPS spectrograph to simultaneously observe a Fabry-Perot lamp and trace possible instrumental drift down to the sub-metre per second level (Wildi et al., 2010). We reduced the data using the dedicated HARPS data reduction software (DRS; Pepe et al., 2002; Lovis, C. & Pepe, F., 2007) and computed the cross-correlation function (CCF) for each spectrum using a G2 numerical mask (Baranne et al., 1996).
We then performed a skew-normal (SN) fit on each CCF (Simola et al., 2019) in orderto extract the stellar radial velocity along with its error, the full width at half maximum (FWHMSN), the contrast (), and the skewness parameter () of the CCF. The advantages of using an SN-fit rather than a Normal fit are thoroughly discussed in Simola et al. (2019), while the SN-fitting details are outlined in, for example, Bonfanti et al. (2023); Luque et al. (2023); Fridlund et al. (2024). After the SN-based extraction, we ended up with an RV time series, whose ancillary vectors (FWHMSN, , ) are the activity indicators used to constrain the polynomial basis to model and de-trend the RV component of the stellar activity (see Table 6).
As stellar activity is not stationary, the correlations between the RV observations and the activity indicators are expected to change over time as discussed in Simola et al. (2022), who proposed to apply the breakpoint (bp) technique (Bai & Perron, 2003) to check whether these correlation changes are statistically significant. If so, the bp algorithm finds the optimal locations of correlation changes, which defines the segmentation characterised by the lowest Bayesian Information Criterion (BIC; Schwarz, 1978). Indeed, we found one breakpoint at observation 48 (BJD = 2458667.940719; with respect to the zero-breakpoint solution). Thus, we split our RV time series in two piecewise stationary segments and de-trended it on a chunk-wise base, rather than performing a global de-trending over the whole time series, similarly to what has already been done in Bonfanti et al. (2023); Luque et al. (2023).
@@ 4 Methods
We jointly analysed the TESS LCs and the RV time series within a Markov chain Monte Carlo (MCMC) framework using the MCMCI code (Bonfanti & Gillon, 2020). When fitting the TESS LCs extracted from Sector 3 that has a cadence of 30 min, we generated the transit model using a cadence of 2 min (the same as the other TESS sectors) and then rebinned it to 30 min following Kipping (2010), who warns that long-cadence photometry may lead to retrieve erroneous system parameters. We imposed Normal priors on the input stellar parameters (i.e. , [Fe/H], , and ), as derived in Sect. 2 with a twofold aim: (i) the induced prior on the mean stellar density (via and ) helps the convergence of the transit model; (ii) limb darkening (LD) coefficients for the TESS filter may be retrieved following interpolation within Atlas9-based666http://kurucz.harvard.edu/grids.html. grids that were pre-computed using the get_lds.py code777https://github.com/nespinoza/limb-darkening by Espinoza & Jordán (2015). We then set Normal priors on the quadratic LD coefficients using the values coming from the grid interpolation (i.e. and ) after re-parameterising them following Holman et al. (2006). On the planetary side, we adopted unbounded (except for the physical limits) uniform priors on the transit depth d, the impact parameter , the orbital period , the transit timing , and the RV semi-amplitude , while we set the eccentricity and the argument of periastron for all planets. We come back to the assumption on the eccentricity below. The specific parameterisations of the jump parameters (aka step parameters) are outlined in Bonfanti & Gillon (2020).
@@
The LCs and the RV time series were de-trended simultaneously during the MCMC analysis using polynomials. To assess the polynomial orders to be associated with the different de-trending parameters for each time series, we first launch several MCMC preliminary runs made of 10 000 steps where we varied only one polynomial order at a time. We then selected the best de-trending baseline (see Table 2) as the one having the lowest BIC. After that, we performed a preliminary MCMCI run comprising 200 000 steps to evaluate the contribution of both the white and red noise in the LCs following Pont et al. (2006); Bonfanti & Gillon (2020), so to properly rescale the photometric errors and get reliable uncertainties on the output parameters. Finally, three independent MCMCI runs comprising 200 000 steps each (burn-in length equal to 20%) were performed to build the posterior distributions of the output parameters after checking their convergence via the Gelman-Rubin statistic (; Gelman & Rubin, 1992).
@@
We also tested the possibility of eccentric orbits by imposing uniform priors on (, ) either bounded to imply or completely unbounded (except for the physical limits). The wider the eccentricity range to be explored by the MCMC scheme, the poorer the parameter convergence, which suggests that the available data are not enough to constrain the planetary eccentricities well. Moreover, the MCMCI runs with are disfavoured by the BIC criterion (e.g. Kass & Raftery, 1995; Trotta, 2007) as well, in fact we obtained . This is also in agreement with the simulations performed by Vanderburg et al. (2019), who suggested that the periods’ commensurability state of planets b and c is more likely maintained if the system is characterised by low eccentricities. Therefore, we adopted the circular solution as the reference one.
5 LC and RV data analysis results
5.1 Joint LC and RV analysis with linear ephemerides